George Young Industries (GYI) acquired industrial robots at the beginning of 2018 and added them to the company’s assembly process. During 2021, management became aware that the $1.3 million cost of the equipment was inadvertently recorded as repair expense on GYI’s books and on its income tax return. The industrial robots have 10-year useful lives and no material salvage value. This class of equipment is depreciated by the straight-line method for financial reporting purposes and for tax purposes it is considered to be MACRS 7-year property. Cost deducted over 7 years by the modified accelerated recovery system as follows:

Year MACRS
Deductions
2018 $ 185,770
2019 318,370
2020 227,370
2021 162,370
2022 116,090
2023 115,960
2024 116,090
2025 57,980
Totals $ 1,300,000


The tax rate is 25% for all years involved.

Required:
1. & 3. Prepare any journal entry necessary as a direct result of the error described and the adjusting entry for 2021 depreciation.
2. Will GYI account for the change (a) retrospectively or (b) prospectively?
Answer #1

2 AB Ans: 1)&3) Year Mars Deduction Straight Line Depreciation Difference Cummulative temporary Differnce Deferred Tax Liabil=E3 Difference =C3-D3 =C4-D4 =C5-D5 =SUME3:E5) Cummulative temporary Differnce Deferred Tax Liability =F3*25% =E3+E4 =F4*25%